# SceneGen: Single-Image 3D Scene Generation in One Feedforward Pass

### v1

@@ Abstract
3D content generation has recently attracted significant research interest due to its applications in VR/AR and embodied AI. In this work, we address the challenging task of synthesizing multiple 3D assets within a single scene image. Concretely, our contributions are fourfold: (i) we present SceneGen, a novel framework that takes a scene image and corresponding object masks as input, simultaneously producing multiple 3D assets with geometry and texture. Notably, SceneGen operates with no need for optimization or asset retrieval; (ii) we introduce a novel feature aggregation module that integrates local and global scene information from visual and geometric encoders within the feature extraction module. Coupled with a position head, this enables the generation of 3D assets and their relative spatial positions in a single feedforward pass; (iii) we demonstrate SceneGen’s direct extensibility to multi-image input scenarios. Despite being trained solely on single-image inputs, our architectural design enables improved generation performance with multi-image inputs; and (iv) extensive quantitative and qualitative evaluations confirm the efficiency and robust generation abilities of our approach. We believe this paradigm offers a novel solution for high-quality 3D content generation, potentially advancing its practical applications in downstream tasks. The code and model will be publicly available at: https://mengmouxu.github.io/SceneGen.
@@
11footnotetext: *: These authors contribute equally to this work.22footnotetext: $\dagger$: Corresponding author.
@@ 1 Introduction
The growing demand for immersive digital environments in applications such as virtual/augmented reality (VR/AR) and embodied AI has spurred significant advancements in 3D content generation [14, 13, 9, 8, 6, 28]. While early efforts primarily focused on synthesizing individual 3D assets [59, 33, 69], recent research focus has shifted to the more challenging task of 3D scene generation. Generating realistic 3D scenes [5, 15, 19, 64, 65, 11, 16], conditioned on input text or images, requires synthesizing multiple assets with accurate geometry, texture, and spatial relationships. This primarily hinges on two key capabilities: (i) 3D asset generation for generating plausible asset geometric topologies from limited textual or visual input. (ii) spatial arrangement for managing inter-object spatial relationships to correctly handle support, occlusion, and other physical interactions among assets.
@@
Existing approaches can be divided into two paradigms: (i) retrieval-based methods [49, 12, 61, 38] typically employ LLMs for layout planning and retrieve matching 3D assets from existing libraries to assemble scenes. While straightforward, their flexibility is limited by the coverage of available assets. (ii) two-stage approaches [18, 35, 62] first generate individual 3D assets, then employ Vision-Language Models (VLMs) or optimization techniques to refine scene structure and spatial arrangement. Although more flexible, their reliance on iterative optimization inevitably leads to inefficiency and error accumulation. The most relevant works to ours are MIDI [23] and PartCrafter [34], which generate multiple assets or parts from a single image. However, they still suffer from limited synthesis fidelity and inaccurate spatial relationships among assets.
@@
To address the aforementioned challenges, we propose SceneGen, a novel 3D scene generation model designed to simultaneously generate multiple assets, including their geometry, texture, and spatial positions, within a scene image, all in a single feedforward pass. Concretely, our framework builds upon an existing single-asset generation model [59] and introduces three key modules, including feature extraction, feature aggregation, and output module. First, the feature extraction module strategically leverages off-the-shelf visual [42] and geometric [53] encoders to extract both asset-level and scene-level features within the scene.
@@
Subsequently, our proposed feature aggregation module, composed of local and global attention blocks, effectively integrates the extracted visual and geometric features while enabling inter-asset interactions during generation to ensure plausible geometric topologies. Benefiting from thorough scene information utilization, the latent features generated by SceneGen can be directly decoded into assets’ relative position, geometry, and texture through the position head and pretrained structure decoder within our output module.
@@
Moreover, despite being trained solely on single-image samples, SceneGen demonstrates remarkable generalization capability to multi-image input scenarios, achieving even better generation quality, which primarily stems from our dedicated architectural design. To conduct a comprehensive and reliable evaluation of SceneGen’s performance on 3D scene generation, we systematically adopt multiple metrics focusing on both geometric and visual quality. Both quantitative and qualitative results demonstrate that our proposed SceneGen significantly outperforms previous methods in terms of generation quality and efficiency, which can generate textured scenes with four assets in about 2 minutes.
@@
The rest of this paper is organized as follows: In Sec. 2, we provide a comprehensive review and discussion of related literature; Sec. 3 elaborates on details of our proposed SceneGen; Sec. 4 presents extensive quantitative and qualitative evaluations; Finally, Sec. 5 concludes with key insights and contributions of our work. To our knowledge, SceneGen is the first 3D scene generation model capable of simultaneously synthesizing geometry, texture, and relative positions of multiple 3D assets in a single feedforward pass, without requiring optimization. We believe this work will inspire advances in high-quality, efficient 3D content generation and facilitate applications in downstream tasks.
@@ 2 Related Work
3D asset synthesis. Typically, 3D asset synthesis aims to generate object-centric geometry and texture from text or image inputs. The recent success of diffusion models [22] in 2D generation [47, 43, 37, 56, 57] has inspired the development of learning-based, scalable 3D content [14, 13, 9, 8, 6, 28] generation, which produce 3D asset in various representations, including explicit forms such as point clouds [40], voxels [41, 25], and SDFs [4, 32], as well as implicit ones like 3D Gaussians [20, 67] and NeRFs [1, 31, 60]. Subsequent advances leverage VAEs [29] for compressing 3D geometry or textures [59, 33, 69] and adopt hybrid mesh-texture pipelines [58, 66, 24, 26], with TRELLIS [59] demonstrating scalable, high-fidelity generation via structured latents. Nevertheless, these methods remain restricted to single-asset synthesis and fundamentally lack the ability to model complex multi-asset scenes.
3D scene generation. Beyond single asset synthesis, 3D scene generation is more challenging yet valuable, aiming to produce multiple coordinated, physically plausible assets within a scene. Prior text-based approaches primarily leverage LLMs for layout planning [49, 12, 61, 38] and retrieve suitable assets from existing libraries. Subsequent image-based methods employ segmentation [5, 15, 19, 64, 10], scene graph [65, 11, 16] and depth/point cloud alignment [50, 62, 10] to assist multi-asset generation and arrangement. As depicted in Figure 2 (b), recent optimization-based methods [18, 35, 62] adopt VLMs for post-processing, refining scene structures through image/text-guided adjustments, but inevitably suffer from inefficiency. Other works (Figure 2 (c)), such as MIDI [23] and PartCrafter [34], explore scene generation conditioned on a single image, but inherently sacrifice reconstruction fidelity due to their canonical-space representations. Our proposed SceneGen uniquely overcomes these limitations by integrating asset-level and scene-level features, enabling robust and efficient 3D scene generation.
@@ 3 Method
In this work, we present SceneGen, designed to jointly perform 3D asset generation within scenes and predict relative spatial positions among assets. Here, we first formally describe our problem formulation in Sec. 3.1; followed by elaboration on our model architecture and training methodology in Sec. 3.2 and 3.3, respectively; finally, we extend SceneGen to multi-view input scenarios in Sec. 3.4.
@@ 3.1 Problem Formulation
$\{(\boldsymbol{S}_{i},\boldsymbol{P}_{i})\}_{i=1}^{N}=\boldsymbol{\mathcal{G}}_{\mathrm{Scene}}(\{\boldsymbol{m}_{i}\}_{i=1}^{N},\boldsymbol{I}_{\mathrm{Scene}})$
@@ 3.2 SceneGen
Our proposed SceneGen ($\boldsymbol{\mathcal{G}}_{\mathrm{Scene}}$) comprises three key stages: (i) feature extraction, employing a scene visual encoder ($\mathrm{\Phi}_{V}$) and a scene geometric encoder ($\mathrm{\Phi}_{G}$) to extract visual and structural features within the scene, implemented using off-the-shelf DINOv2 [42] and VGGT [53], respectively; (ii) feature aggregation, comprising $M$ DiT [43] blocks, each integrating a local attention block, a global attention block, and a feedforward network; and (iii) output module, which introduces a position head ($\Psi_{\mathrm{pos}}$) for predicting the spatial locations of assets and adopts off-the-shelf sparse structure (SS) and structured latents (SLAT) decoders [59] for decoding scene geometry structures. By integrating these complementary modules, our SceneGen effectively captures both local asset-level and global scene-level features, enabling it to simultaneously generate multiple 3D assets and predict their relative positions.
Feature extraction. Our SceneGen starts with extracting both local and global features from a given scene image ($\boldsymbol{I}_{\mathrm{scene}}$) with the visual encoder ($\Phi_{V}$) and geometric encoder ($\Phi_{G}$). Specifically, for each object with its corresponding segmentation mask ($\boldsymbol{m}_{i}$), we obtain four complementary feature representations: (i) the object’s individual visual features ($\mathcal{F}_{i}^{V}$); (ii) the visual features of its mask ($\mathcal{F}_{i}^{\mathrm{mask}}$); (iii) scene global visual features ($\mathcal{F}_{\mathrm{global}}^{V}$); and (iv) the global geometric features ($\mathcal{F}_{\mathrm{global}}^{\mathrm{{geo}}}$), denoted as:
@@
Here, $\otimes$ represents pixel-wise multiplication. These features are then concatenated along the sequence dimension into a unified scene context ($\mathcal{F}_{i}^{\mathrm{scene}}$), formulated as:
@@
Feature aggregation. Subsequently, SceneGen employs a feature aggregation module to integrate the extracted scene context features ($\mathcal{F}_{i}^{\mathrm{scene}}$), enabling simultaneous generation of multiple 3D assets. The module consists of a local attention block that refines details of individual assets, a global attention block that incorporates scene context information and facilitates inter-asset interactions during generation, and a feedforward network. Specifically, we parameterize the local attention blocks and feedforward networks with pre-trained weights from TRELLIS [59], a flow-matching [36] model that sytnehsizes 3D content from noisy sparse structure latents. For clarity and conciseness, given the sparse structure latents ($\{\mathbf{x}_{i}\}_{i=1}^{N}$, where each ${\mathbf{x}}_{i}\in\mathbb{R}^{T\times C}$) of $N$ objects within a scene, we denote the standard attention mechanism as $\mathrm{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})$, and elaborate on a single DiT block as follows.
The local attention block aims to enhance details of individual objects through asset-level self-attention (AS) and cross-attention (AC). Concretely, it focuses on fusing the latent features of each object ($\mathbf{x}_{i}$) with their corresponding visual features ($\mathcal{F}_{i}^{V}$) to yield refined representations of each object ($\mathbf{x}_{i}^{\mathrm{AC}}$), which can be formulated as:
@@
To establish inter-dependencies among 3D assets, we propose the global attention block, comprising scene-level self-attention (SS) and cross-attention (SC), which capture inter-object relationships and integrate scene geometry, respectively. This ensures physically plausible spatial arrangements of generated assets.
@@
Output module. After passing through $M$ DiT blocks, we obtain the updated position tokens ($\{\hat{\boldsymbol{p}}\}_{i=1}^{N}$) and latent features ($\{\tilde{\mathbf{x}}\}_{i=1}^{N}$) of each generated asset, which are then decoded into their relative spatial positions ($\{\boldsymbol{P}\}_{i=1}^{N}$), and structural and texture representations ($\{\boldsymbol{S}\}_{i=1}^{N}$), respectively. For relative positions, we extract and concatenate the position tokens of all non-query assets, which are then decoded into corresponding 8D position vectors ($\{\hat{\boldsymbol{P}}_{i}\}_{i=2}^{N}$) by our proposed position head ($\Psi_{\mathrm{pos}}$), comprising four self-attention layers and a linear layer, denoted as:
@@
Here, each spatial position vector ($\hat{\boldsymbol{P}}_{i}$) represents an asset’s spatial position (translation, rotation, and scale) relative to the pre-selected query asset ($\boldsymbol{q}_{\mathrm{query}}$). Additionally, the latent features can be directly decoded into the geometry and texture of each asset ($\{\boldsymbol{S}\}_{i=1}^{N}$) using off-the-shelf sparse structure generator ($\boldsymbol{\mathcal{G}}_{\mathrm{S}}$) and structured latents generator ($\boldsymbol{\mathcal{G}}_{\mathrm{L}}$) in TRELLIS [59], represented as:
$\{\boldsymbol{S}\}_{i=1}^{N}=\boldsymbol{\mathcal{G}}_{\mathrm{L}}(\boldsymbol{\mathcal{G}}_{\mathrm{S}}(\{\tilde{\mathbf{x}}\}_{i=1}^{N}))$
@@ 3.3 Training
Training data. Our SceneGen model is trained on 3D-FUTURE [14], containing photorealistic scene renderings with instance masks and asset annotations. This dataset comprises 12K training scenes and 4.8K test scenes, each featuring a scene image with one or multiple objects. To better capture inter-object spatial relationships, we augment the training set by iteratively designating each asset as the query asset, and randomly permuting the remaining assets, which expands the effective training samples to 30K.
Training objectives. Our SceneGen model is trained end-to-end using a composite loss function ($\mathcal{L}$) comprising three key components: (i) the average conditional flow matching [36] loss ($\mathcal{L}_{\mathrm{cfm}}$), on each generated asset for supervising asset generation; (ii) the position loss ($\mathcal{L}_{\mathrm{pos}}$) for maintaining accurate relative spatial arrangements among assets; and (iii) the voxel-space collision loss ($\mathcal{L}_{\mathrm{coll}}$) for enforcing physically plausible object placements. The overall objective function ($\mathcal{L}$) combines these components with a weighting factor ($\lambda$), which can be formulated as:
@@
Concretely, the flow matching loss establishes straight probability paths between distributions via linear interpolation: $\mathbf{x}_{i}(t)=(1-t)\mathbf{x}_{i}^{0}+t\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, $t\in[0,1]$, and $\mathbf{x}_{i}^{0}$ denotes the noise-free sparse structure latents for each of the $N$ assets in the scene. The conditional flow matching objective learns a parameterized function $\boldsymbol{v}_{\theta}$ to approximate the velocity field ($\boldsymbol{v}(\mathbf{x}_{i}(t),t)=\nabla_{t}\mathbf{x}_{i}(t)$), represented as:
@@
The collision loss ($\mathcal{L}_{\mathrm{coll}}$) quantifies surface collision in a $64\times 64\times 64$ voxel grid ($\boldsymbol{V}$). Concretely, the predicted sparse structure latents ($\tilde{\mathbf{x}}_{i}$) are decoded into point clouds ($\{\boldsymbol{p}_{i}\}_{i=1}^{L}$) via a pretrained sparse structure decoder from TRELLIS [59], then transformed using predicted pose parameters ($\hat{\boldsymbol{P}}_{i}$) and voxelized into $\boldsymbol{V}$. The collision loss is defined as the ratio of overlapping surface voxels to all surface voxels, using the Huber loss ($\|\cdot\|_{\delta_{C}}$), denoted as:
@@
Here, $\mathrm{IoU}_{\mathrm{scene}}=0$ means there are no asset collisions.

### v2

@@ Abstract
3D content generation has recently attracted significant research interest, driven by its critical applications in VR/AR and embodied AI. In this work, we tackle the challenging task of synthesizing multiple 3D assets within a single scene image. Concretely, our contributions are fourfold: (i) we present SceneGen, a novel framework that takes a scene image and corresponding object masks as input, simultaneously producing multiple 3D assets with geometry and texture. Notably, SceneGen operates with no need for extra optimization or asset retrieval; (ii) we introduce a novel feature aggregation module that integrates local and global scene information from visual and geometric encoders within the feature extraction module. Coupled with a position head, this enables the generation of 3D assets and their relative spatial positions in a single feedforward pass; (iii) we demonstrate SceneGen’s direct extensibility to multi-image input scenarios. Despite being trained solely on single-image inputs, our architecture yields improved generation performance when multiple images are provided; and (iv) extensive quantitative and qualitative evaluations confirm the efficiency and robustness of our approach. We believe this paradigm offers a novel solution for high-quality 3D content generation, potentially advancing its practical applications in downstream tasks. The code and model will be publicly available at: https://mengmouxu.github.io/SceneGen.
@@
11footnotetext: *: These authors contribute equally to this work.
@@ 1 Introduction
The growing demand for immersive digital environments in applications such as virtual/augmented reality (VR/AR) and embodied AI has spurred significant advancements in 3D content generation [14, 13, 9, 8, 6, 28]. While early efforts primarily focus on synthesizing individual 3D assets [59, 33, 69], recent research attention has shifted to the more challenging task of 3D scene generation. Generating realistic 3D scenes [5, 15, 19, 64, 65, 11, 16], whether conditioned on input text or images, requires synthesizing multiple assets with accurate geometry, texture, and spatial relationships. This challenge fundamentally hinges on two key capabilities: (i) 3D asset generation for creating plausible asset geometric topologies from limited textual or visual input; and (ii) spatial arrangement for managing inter-object spatial relationships to ensure physical plausibility, such as support, occlusion, and other interactions among assets.
@@
In general, existing works fall into two paradigms: (i) retrieval-based methods [49, 12, 61, 38] typically employ LLMs for layout planning and retrieve matching 3D assets from existing libraries to assemble scenes. Though straightforward, their flexibility is constrained by the coverage of available asset libraries; (ii) two-stage approaches [18, 35, 62] first synthesize individual 3D assets and then employ vision-language models (VLMs) or optimization techniques to refine scene structure and spatial arrangement. While more flexible, their reliance on iterative optimization inevitably leads to inefficiency and error accumulation. The most relevant works to ours are PartCrafter [34] and MIDI [23], which generate parts or multiple assets from a single image. However, they still suffer from limited synthesis fidelity and inaccurate spatial relations among assets.
@@
To tackle the aforementioned challenges, we propose SceneGen, a novel 3D scene generation model designed to simultaneously generate multiple assets, including their geometry, texture, and spatial positions, from a single scene image in a single feedforward pass (Figure 1). Concretely, our framework builds upon an existing single-asset generation model [59] and incorporates three key modules: feature extraction, feature aggregation, and output.
@@
Specifically, the feature extraction module first strategically leverages off-the-shelf visual [42] and geometric [53] encoders to extract both asset-level and scene-level representations. Subsequently, our proposed feature aggregation module, composed of local and global attention blocks, effectively integrates these visual and geometric features while facilitating inter-asset interactions during generation to ensure geometrically plausible topologies. Finally, leveraging this comprehensive scene context, the output module can directly decode the generated latent features into the assets’ relative position, geometry, and texture via a position head and a pre-trained structure decoder.
@@
Moreover, despite being trained exclusively on single-image samples, SceneGen exhibits remarkable generalization to multi-image input scenarios, yielding even better generation quality, which primarily stems from our dedicated architectural design. To ensure a comprehensive and reliable evaluation of SceneGen, we systematically adopt multiple metrics focusing on both geometric and visual quality. Both quantitative and qualitative results demonstrate that our proposed SceneGen significantly outperforms previous methods in terms of generation quality and efficiency, which can generate a textured scene with four assets in approximately 2 minutes on a single A100 GPU.
@@
The rest of this paper is organized as follows: Sec. 2 provides a comprehensive review and discussion of related literature. Sec. 3 elaborates on our proposed SceneGen framework. Sec. 4 presents extensive quantitative and qualitative evaluations. Finally, Sec. 5 concludes with key insights and contributions. To our knowledge, SceneGen is the first 3D scene generation model capable of simultaneously synthesizing geometry, texture, and relative positions of multiple 3D assets in a single feedforward pass, without requiring per-scene optimization. We believe this work will inspire future advances in high-quality, efficient 3D content generation and facilitate diverse downstream applications.
@@ 2 Related Work
3D asset synthesis. Typically, 3D asset synthesis aims to generate object-centric geometry and texture from text or image inputs. The recent success of diffusion models [22] in 2D generation [47, 43, 37, 56, 57] has inspired the development of learning-based, scalable 3D content [14, 13, 9, 8, 6, 28] generation, which produce 3D asset in various representations, including explicit forms such as point clouds [40], voxels [41, 25], and SDFs [4, 32], as well as implicit ones like 3D Gaussians [20, 67] and NeRFs [1, 31, 60]. Subsequent advances leverage VAEs [29] for compressing 3D geometry or textures [59, 33, 69] and adopt hybrid mesh-texture pipelines [58, 66, 24, 26], with TRELLIS [59] demonstrating scalable, high-fidelity generation via structured latents. Nevertheless, these methods remain restricted to single-asset synthesis and fundamentally lack the capability to model complex multi-asset scenes.
3D scene generation. Beyond single-asset synthesis, 3D scene generation is more challenging yet valuable, aiming to produce multiple coordinated, physically plausible assets within a scene. Prior text-based approaches primarily leverage LLMs for layout planning [49, 12, 61, 38] and retrieve suitable assets from existing libraries. Subsequent image-based methods employ segmentation [5, 15, 19, 64, 10], scene graphs [65, 11, 16] and depth/point cloud alignment [50, 62, 10] to assist in multi-asset generation and arrangement. As depicted in Figure 2 (b), recent optimization-based methods [18, 35, 62] adopt VLMs for post-processing, refining scene structures via image- or text-guided adjustments, but inevitably suffer from inefficiency. Other works (Figure 2 (c)), such as MIDI [23] and PartCrafter [34], explore scene generation conditioned on a single image, but inherently sacrifice reconstruction fidelity due to their reliance on canonical-space representations. To overcome these limitations, our proposed SceneGen uniquely integrates asset-level and scene-level features, enabling robust and efficient 3D scene generation.
@@ 3 Method
In this work, we present SceneGen, designed to jointly perform 3D asset generation within scenes and predict relative spatial positions among assets. Here, we first formally describe our problem formulation in Sec. 3.1; followed by elaboration on our model architecture and training method in Sec. 3.2 and Sec. 3.3, respectively; finally, we extend SceneGen to multi-view input scenarios in Sec. 3.4.
@@ 3.1 Problem Formulation
$\{(\boldsymbol{S}_{i},\boldsymbol{P}_{i})\}_{i=1}^{N}=\boldsymbol{\mathcal{G}}_{\mathrm{Scene}}(\boldsymbol{I}_{\mathrm{Scene}},\{\boldsymbol{m}_{i}\}_{i=1}^{N})$
@@ 3.2 SceneGen
SceneGen framework ($\boldsymbol{\mathcal{G}}_{\mathrm{Scene}}$) comprises three key stages: (i) feature extraction, employing a scene visual encoder ($\mathrm{\Phi}_{V}$) and a scene geometric encoder ($\mathrm{\Phi}_{G}$) to extract visual and structural features within the scene, implemented using pre-trained DINOv2 [42] and VGGT [53], respectively; (ii) feature aggregation, comprising $M$ DiT [43] blocks, each integrating a local attention block, a global attention block, and a feedforward network; and (iii) output module, which introduces a position head ($\Psi_{\mathrm{pos}}$) for predicting the spatial locations of assets and adopts off-the-shelf sparse structure (SS) and structured latents (SLAT) decoders [59] for decoding scene geometry structures. By integrating these complementary modules, our SceneGen effectively captures both local asset-level and global scene-level features, enabling it to simultaneously generate multiple 3D assets and predict their relative positions.
Feature extraction. SceneGen starts with extracting both local and global features from a given scene image ($\boldsymbol{I}_{\mathrm{scene}}$) with the visual encoder ($\Phi_{V}$) and geometric encoder ($\Phi_{G}$). Specifically, for each object with its corresponding segmentation mask ($\boldsymbol{m}_{i}$), we obtain four complementary feature representations: (i) the object’s individual visual features ($\mathcal{F}_{i}^{V}$); (ii) the visual features of its mask ($\mathcal{F}_{i}^{\mathrm{mask}}$); (iii) scene global visual features ($\mathcal{F}_{\mathrm{global}}^{V}$); and (iv) the global geometric features ($\mathcal{F}_{\mathrm{global}}^{\mathrm{{geo}}}$), formulated as:
@@
Here, $\otimes$ denotes pixel-wise multiplication. These features are then concatenated along the sequence length dimension into a unified scene context ($\mathcal{F}_{i}^{\mathrm{scene}}$), formulated as:
@@
Feature aggregation. To integrate the extracted scene context features ($\mathcal{F}_{i}^{\mathrm{scene}}$), SceneGen employs a feature aggregation module, that enables the simultaneous generation of multiple 3D assets. This module comprises a local attention block that refines details of individual assets, a global attention block that incorporates scene context to facilitate inter-asset interactions, and a feedforward network. Concretely, the local attention blocks and feedforward networks are initialised from pre-trained weights from TRELLIS [59], a flow-matching [36] model designed to synthesize 3D content from noisy sparse structure latents. For clarity and conciseness, given the sparse structure latents ($\{\mathbf{x}_{i}\}_{i=1}^{N}$, where each ${\mathbf{x}}_{i}\in\mathbb{R}^{T\times C}$) of $N$ objects in a scene, we denote the standard attention mechanism as $\mathrm{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})$, and elaborate on a single DiT block as follows.
The local attention block aims to enhance details of individual objects through asset-level self-attention (AS) and cross-attention (AC). To be specific, it focuses on fusing the latent features of each object ($\mathbf{x}_{i}$) with their corresponding visual features ($\mathcal{F}_{i}^{V}$) to yield refined representations of each object ($\mathbf{x}_{i}^{\mathrm{AC}}$), which can be formulated as:
@@
To establish inter-dependencies among 3D assets, we propose a global attention block, comprising scene-level self-attention (SS) and cross-attention (SC), which capture inter-object relationships and integrate scene geometry, respectively. Consequently, this design ensures physically plausible spatial arrangements of generated assets.
@@
Output module. After passing through $M$ DiT blocks, we obtain the updated position tokens ($\{\hat{\boldsymbol{p}}\}_{i=1}^{N}$) and latent features ($\{\tilde{\mathbf{x}}\}_{i=1}^{N}$) of each generated asset, which are subsequently decoded into their relative spatial positions and detailed 3D representations (structure and texture), respectively. For relative positions, we concatenate the position tokens of all non-query assets, which are then decoded into corresponding 8D position vectors ($\{\hat{\boldsymbol{P}}_{i}\}_{i=2}^{N}$) by our proposed position head ($\Psi_{\mathrm{pos}}$), comprising four self-attention layers and a linear layer, denoted as:
@@
Here, each vector ($\hat{\boldsymbol{P}}_{i}$) represents an asset’s spatial position (translation, rotation, and scale) relative to the pre-selected query asset ($i=1$). Furthermore, the latent features can be directly decoded into the geometry and texture of each asset ($\{\hat{\boldsymbol{S}}\}_{i=1}^{N}$) using off-the-shelf sparse structure generator ($\boldsymbol{\mathcal{G}}_{\mathrm{S}}$) and structured latents generator ($\boldsymbol{\mathcal{G}}_{\mathrm{L}}$) from TRELLIS [59], formulated as:
$\{\hat{\boldsymbol{S}}\}_{i=1}^{N}=\boldsymbol{\mathcal{G}}_{\mathrm{L}}(\boldsymbol{\mathcal{G}}_{\mathrm{S}}(\{\tilde{\mathbf{x}}\}_{i=1}^{N}))$
@@ 3.3 Training
Training data. Our SceneGen model is trained on the 3D-FUTURE [14] dataset, containing photorealistic scene renderings with instance masks and asset annotations. This dataset comprises 12K training scenes and 4.8K test scenes, each featuring a scene image with one or multiple objects. To better capture inter-object spatial relationships, we augment the training set by iteratively designating each asset as the query asset while randomly permuting the remaining assets, which expands the effective training samples to 30K.
Training objectives. Our SceneGen model is trained end-to-end using a composite loss function ($\mathcal{L}$) comprising three key components: (i) the average conditional flow matching [36] loss ($\mathcal{L}_{\mathrm{cfm}}$), applied to each generated asset for supervising asset generation; (ii) the position loss ($\mathcal{L}_{\mathrm{pos}}$) for maintaining accurate relative spatial arrangements among assets; and (iii) the voxel-space collision loss ($\mathcal{L}_{\mathrm{coll}}$) for enforcing physically plausible object placement. The overall objective function ($\mathcal{L}$) combines these components with a weighting factor ($\lambda$), which can be formulated as:
@@
Concretely, the flow matching loss establishes straight probability paths between distributions via linear interpolation: $\mathbf{x}_{i}(t)=(1-t)\mathbf{x}_{i}^{0}+t\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, $t\in[0,1]$, and $\mathbf{x}_{i}^{0}$ denotes the noise-free sparse structure latents for each of the $N$ assets. The conditional flow matching objective ($\mathcal{L}_{\mathrm{cfm}}$) learns a parameterized function $\boldsymbol{v}_{\theta}$ to approximate the velocity field ($\boldsymbol{v}(\mathbf{x}_{i}(t),t)=\nabla_{t}\mathbf{x}_{i}(t)$), represented as:
@@
The collision loss ($\mathcal{L}_{\mathrm{coll}}$) quantifies surface collision in a $64\times 64\times 64$ voxel grid ($\boldsymbol{V}$). Specifically, the predicted sparse structure latents ($\tilde{\mathbf{x}}_{i}$) are decoded into point clouds ($\{\boldsymbol{p}_{i}\}_{i=1}^{L}$) via a pre-trained sparse structure decoder from TRELLIS [59], then transformed using predicted pose parameters ($\hat{\boldsymbol{P}}_{i}$) and voxelized into $\boldsymbol{V}$. The collision loss is defined as the ratio of overlapping surface voxels to all surface voxels, using the Huber loss ($\|\cdot\|_{\delta_{C}}$), denoted as:
@@
Ideally, $\mathrm{IoU}_{\mathrm{scene}}=0$ indicates there are no asset collisions.